OpenStudy (anonymous):

in a 2003 study, the Accreditation Council for Graduate Medical Education found that medical residents' mean number of hors worked in a week in 81.7. suppose the number of hours worked per week by medical residents is approximately normally distributed with a standard deviation of 6.9 hours. 1) determine the 75th percentile for the number of hours worked in a week by medical residents. 2) determine the number of hours worked in a week that makes up the middle 80% of medical residents. justify and explain your reasoning

OpenStudy (anonymous):

You need to convert the figure of 75 hours to a z-value (because this is the parameter used in the normal distribution tables), using the formula z = (x - μ )/s.d So z = - 6.7 / s.d. But when you are considering the mean of a sample (as opposed to the value for a single individual), the standard deviation diminishes as the size of the sample increases, according to the function s.d = σ / √n (where σ is the population standard deviation) So for the first sample, s.d. = 6.9/√5 = 3.0858, and z = - 6.7/3.0858 = - 2.171 And from the table of areas of the normal distribution, P(x < - 2.171) = 0.0148 Second sample similarly, s.d = 6.9/√8 = 2.4395, hence z = - 2.7464, and P(x< - 2.7464) = 0.00301

OpenStudy (anonymous):

Its Right ! (:

OpenStudy (anonymous):

thank you so much for your help guys :) @Yumira

OpenStudy (anonymous):

Welcome :D